# Chapter 1      Introduction to IoT

### 09 Dec 2023 v1

@@ Abstract
The Internet of Things (IoT) has rapidly transformed the 21st century, enhancing decision-making processes and introducing innovative consumer services such as pay-as-you-use models. This integration of smart devices and automation technologies has revolutionized our lives. However, it is essential to recognize the significant concerns surrounding security, privacy, intellectual property rights, safety, and trust in this technological landscape, which require further exploration. This chapter serves as a comprehensive guide to newcomers interested in the IoT domain, providing a foundation for making future contributions. It begins by explaining the core concept of IoT, its historical evolution, network components, and key characteristics. The advantages of IoT deployment across various domains are discussed, along with an introduction to the foundational layered architectures supporting IoT. The chapter also delves into the taxonomy of IoT technologies, encompassing hardware, software, wireless communication technologies, hardware platforms, and cloud solutions. Existing applications in major domains like smart cities, healthcare, and agriculture are explored, offering a broad perspective on the IoT’s application potential. In addressing prevalent issues and challenges in designing and deploying IoT applications, the chapter examines security threats across architectural layers, ethical considerations, user privacy concerns, and trust-related issues. This discussion equips researchers with a solid understanding of diverse IoT aspects, providing a strong foundation for both practical IoT projects and the development of novel theoretical approaches within the IoT field.
Keywords: Taxonomy, Architectural Layers, Ethical Considerations, Security Threats, Hardware Platforms, Trust-related Issues.
@@ 1.1 Introduction
The Internet of Things (IoT) is a network of interconnected devices with sensors, actuators, processors, and various communication technologies. Sensors collect real-time data from both internal states and external surroundings, ranging from mobile phones to microwave ovens. Actuators, in turn, respond to data or commands, enabling automation and remote control of physical devices. Data collected by sensors undergoes processing either at the network edge or on central servers, with some preprocessing occurring directly in the sensors or end devices. Processed data is then transmitted to remote servers for further analysis, storage, and processing. These data form the basis for analysis, decision-making, and subsequent actions, which can be physical (e.g., adjusting a smart thermostat) or virtual (e.g., sending notifications) [4]. The applications of IoT are extensive and diverse, impacting various aspects of our lives. From personal convenience in smart homes to healthcare and fitness innovations, IoT has the potential to influence personal, financial, physical, educational, professional, and mental aspects of individuals. In smart homes, IoT enables remote control of electrical appliances, lighting, coffee brewing, thermostat adjustments, and even hands-free operations through voice commands [5]. In healthcare, wearable IoT devices offer remote monitoring, allowing caregivers and healthcare professionals to provide timely assistance in emergencies. Additionally, individuals can use wearable devices to track sleep patterns, physical activity, and overall fitness [6]. These examples only scratch the surface of IoT’s broad application landscape, indicating the exciting possibilities and challenges that researchers are exploring for the future.
@@
The Internet of Things (IoT) has the potential to transform how people interact with technology, providing greater convenience, efficiency, and personalization in daily life. Despite its transformative impact, IoT faces challenges. The sheer number of devices and the substantial data generated pose significant challenges, with a projected 41.6 billion IoT devices producing 79.4 zettabytes of data by 2025 [7]. Addressing this requires scalable architectures and enhanced processing capabilities. Moreover, IoT heavily relies on wireless communication, leading to challenges such as distortion and unreliability in geographically dispersed locations. Ensuring dependable data transmission becomes a pivotal challenge, emphasizing the critical role of communication technologies in the IoT landscape. Beyond technical hurdles, various general and domain-specific challenges are crucial for the success of IoT. Identifying and addressing these multifaceted challenges collectively is essential to unlock the full potential of IoT and overcome obstacles for broader adoption and integration into our lives.
@@
Motivation of this chapter: IoT is not a novel concept; researchers have been exploring this field for decades. Consequently, the question arises: what sets this chapter apart from others, and why should readers invest their time in exploring the basics presented here? While there are several book chapters introducing IoT, many share outdated concepts, lacking the latest insights from the ongoing exploration in this dynamic field. For instance, in [8], only IoT applications are discussed, and in [9], authors overlook major domains, including IoT applications, advantages, challenges, and technologies. Notably, Nagaraj et al. [10] present a well-structured discussion covering technologies, architectures, applications, and challenges. However, their discussion is limited to these aspects, with a less extensive exploration of the application section. Therefore, the need arises for a comprehensive chapter that covers all aspects and incorporates recent additions in the IoT field. This chapter goes beyond the basics, encompassing fundamental components, characteristics, and advantages. It delves into architectures, provides a taxonomy of technologies used in IoT, explores a significant number of applications across diverse domains, and addresses ethical considerations as well as legal and regulatory issues. In essence, our chapter serves as a holistic guide, covering the most significant facets of IoT, from foundational principles to emerging research challenges and future directions. Readers are encouraged to explore this chapter for a thorough and up-to-date understanding of the evolving landscape of IoT.
Authors’ Contributions: The authors’ contributions in this chapter can be summarized as follows:
@@
Chapter organization: It begins with an exploration of the fundamentals in Section 1.2, covering the history 1.2.1, components 1.2.2, and characteristics 1.2.3. In Section 1.3, the advantages of IoT are examined, with a focus on both the general perspective and specific domains such as healthcare and agriculture. By introducing a generic architecture in Section 1.4, the taxonomy of technologies that underpin IoT operations are discussed in detail in Section 1.5. Furthermore, Section 1.6 explores the extensive array of application domains of IoT, with a focus on multiple promising research works within each domain, aiming to provide a comprehensive understanding of its real-world applications. It concludes in Section 1.7 by addressing the key challenges that IoT faces today, considering insights from the reviewed literature, and sheds light on potential future scopes and developments in the ever-evolving realm of IoT technology.
@@ 1.2 Fundamentals of IoT
At its core, the IoT revolves around the concept of connecting everyday objects and devices to the internet, enabling them to communicate, collect data, and perform actions autonomously or in response to commands. This basic concept involves three key elements: sensors and actuators that gather and interact with data, a network infrastructure for data transmission, and cloud-based platforms for data storage, processing, and analysis. By interconnecting these elements, IoT empowers us to enhance efficiency, gain real-time insights, and create smart, responsive systems that impact various aspects of our lives, from smart homes and cities to industries and healthcare. The evolution of IoT, which has unfolded over several decades, began with the concept of smart objects. Since then, IoT has undergone numerous groundbreaking transformations that have left the world astounded, thanks to its unique and convenient characteristics. The advantages brought by this technology are unquantifiable, as it touches nearly every aspect of people’s lives, simplifying and improving them. This section comprehensively discusses these fundamental IoT concepts, covering the following topics: The History of IoT, Components of IoT, Characteristics of IoT.
1.2.1 Historical Development of IoT
@@
The 1980s saw a growing interest in enhancing physical objects with sensors and intelligence. Commercial service providers began supporting public access to ARPANET, an early precursor to the modern Internet, during this time. Satellites and landlines played pivotal roles in establishing the foundational communication infrastructure for the emerging IoT. The concept of a network of smart devices was initially explored as early as 1982 when a Coca-Cola vending machine [^3] The ”Only” Coke Machine on the Internet. Available at: Carnegie Mellon University at Carnegie Mellon University was modified to connect to ARPANET. This allowed local programmers to remotely monitor the vending machine’s contents, ensuring drinks were available and cold before making a purchase. However, the technology was challenging to manage, and progress in this field was limited during that period [13]. In parallel, during the 1980s, local area networks (LANs) gained popularity and proved effective for real-time communication and document sharing among groups of PCs.
@@
In the 1990s, advancements in wireless technology set the stage for the widespread adoption of ”machine-to-machine” (M2M) solutions in enterprise and industrial contexts, particularly for equipment monitoring and operation. However, many of these early M2M solutions relied on closed, purpose-built networks and proprietary or industry-specific standards, rather than utilizing Internet Protocol (IP)-based networks and Internet standards [14]. By the mid-1990s, the Internet had expanded its global reach, offering new possibilities for researchers and technologists to explore ways to enhance connections between humans and machines. A significant milestone in this journey was the creation of the first Internet-connected ’device’ by John Romkey—an IP-enabled toaster that could be controlled over the Internet. This innovative toaster was showcased at an Internet conference in 1990, marking an early example of the IoT in action [15].
@@
In 1991, Mark Weiser’s paper “The Computer of the 21st Century” [16] and academic events like UbiComp and PerCom shaped the contemporary IoT vision [14]. Global Positioning Satellites (GPS) became a reality in early 1993 with the Department of Defense establishing a stable system of 24 satellites. Privately owned commercial satellites soon followed, enhancing IIoT functionality [3]. In early 1994, Reza Raji introduced the IoT concept in IEEE Spectrum, describing it as “moving small data packets to integrate and automate from home appliances to entire factories” [17]. Steve Mann’s near-real-time WearCam, powered by a 64-processor setup, was invented later that year. Between 1993 and 1997, companies proposed IoT solutions, including Microsoft’s “at Work” and Novell’s NEST. Momentum grew as Bill Joy introduced device-to-device communication in his ”Six Web” framework at the 1999 World Economic Forum [3].
@@
The term “Internet of Things” was coined by Peter T. Lewis in a 1985 speech during the Congressional Black Caucus Foundation’s 15th Annual Legislative Weekend in Washington, D.C. Lewis defined IoT as the integration of people, processes, and technology with connectable devices and sensors for remote monitoring, status assessment, manipulation, and trend evaluation related to these devices [3]. In 1997, Paul Saffo described sensors and their future role. British technologist Kevin Ashton, while serving as the executive director of the Auto-ID Centre at MIT, independently coined the term “Internet of Things”. During his time at Procter & Gamble, he explored radio-frequency identification (RFID), a technology framework enabling physical devices to connect via microchips and wireless signals. In the same year, they developed a global RFID-based item identification system [14].
@@
In 1999, Kevin Ashton was the first to describe the IoT and proposed the name “Internet of Things” during a presentation for Procter & Gamble. He believed RFID technology, primarily designed for inventory tracking, was a significant prerequisite for the Internet of Things, allowing computers to efficiently manage and monitor individual objects. The concept of tagging objects has been realized through technologies like digital watermarking, barcodes, and QR codes, used for identification and tracking purposes [18]. Subsequent technological advancements, including the proliferation of smartphones, cloud computing, improved processing power, and enhanced software algorithms, along with the availability of sophisticated sensors capable of measuring various parameters, laid the foundation for robust data collection, storage, and processing for the IoT’s growth.
@@
As a significant step forward in commercialising IoT, LG announced plans to launch a smart refrigerator capable of autonomously managing its contents in 2000. Walmart and the US Department of Defense pioneered inventory tracking using RFID and the IoT in $2002-2003$. RFID gained prominence in the US Army’s Savi program in 2003, and Walmart expanded its RFID usage worldwide that same year. In 2004, Cornelius “Pete” Peterson, CEO of NetSilicon, predicted that IoT devices would dominate the next era of information technology, particularly in fields like medical devices and industrial controls [3]. In 2005, numerous articles in mainstream newspapers such as The Guardian, Scientific American, and The Boston Globe discussed IoT’s future direction.
@@
1.2.2 Components of IoT
The IoT consists of several key components that serve as the essential building blocks for constructing an IoT system. This section provides an in-depth exploration of the principal components within IoT. IoT comprises three main components: (1) Sensors/Devices and Actuators, (2) Storage and Data Analytics, and (3) Interpretation and Visualization Tools. Each of these is further categorized into various subcomponents.
@@ 1.2.2 Components of IoT
    Sensors/Devices
    Sensors play a crucial and essential role within an IoT system. Given that IoT operates by gathering data from the surrounding environment, it is necessary for all IoT applications to incorporate one or more sensors to meet this need. A defining characteristic of IoT devices is their context awareness, which is made possible through the utilization of sensor technology. Sensors are not only compact and cost-effective but also energy-efficient. However, they are subject to limitations such as battery capacity and ease of deployment. [4]. An overview of various types of sensors has been provided below.
@@
      Mobile based Sensors
      Smartphones, which are widespread and commonly used, are equipped with various sensors, including accelerometers and environmental monitoring sensors. Given their extensive usage, researchers are exploring the potential of using smartphones as integral components in building smart IoT solutions. These applications can harness sensor data from smartphones to generate valuable insights and outcomes. Some of the general sensors found in smartphones include [19]:
      1. i.
        Accelerometer: This sensor monitors and detects the motion or acceleration of the device, typically measuring changes in the smartphone’s velocity in three dimensions [20]. The data patterns recorded by the accelerometer can be employed to identify the user’s physical activities, including running, walking, and cycling.
      2. ii.
        Gyroscope: The gyroscope introduces an extra dimension compared to the accelerometer by being able to detect rotation or tilting of the phone.
      3. iii.
        GPS: The Global Positioning System (GPS) employs satellites to determine the location of a mobile phone.
      4. iv.
        Magnetometer: The magnetometer detects the magnetic field and can function as a compass.
      5. v.
        Light Sensor: The light sensor measures the intensity of light in the environment around it. It can be applied in different ways, such as managing the room’s lighting or adjusting the screen’s brightness.
      6. vi.
        Proximity Sensor: The proximity sensor employs an infrared (IR) LED that emits infrared rays. These rays bounce off objects they encounter, and by measuring the time it takes for them to return, one can calculate the distance to those objects. This technology enables us to measure how far things are from the phone.
      Certain smartphones, like the Samsung Galaxy S4, come equipped with extra sensors, including a thermometer, barometer, and humidity sensor. These sensors are used to measure temperature, atmospheric pressure, and humidity levels, respectively [4].
@@
      The healthcare industry is one of the most influenced fields where innovation and IoT have paved the way. Wearable devices and sensors have facilitated remote monitoring for physicians and enabled researchers to collect data continuously and in real-time. A variety of wearable devices are now on the market for measuring physiological data, including blood pressure, heart rate, respiration, and glucose levels, [21]. These devices come in different forms, such as wristbands, smartwatches, and monitoring patches. Smartwatches and fitness trackers, known for their versatility, have gained popularity among consumers. Likewise, monitoring patches have emerged as a valuable asset to the healthcare sector by enabling remote treatment for patients.
@@
      Neural sensors play a crucial role in comprehending the workings of human neurons. They enable us to decode brain signals, assess the brain’s current state, and, when necessary, optimize it for improved focus and attention. This practice is commonly referred to as Neurofeedback [22].
@@
      Environmental sensors serve the purpose of measuring various environmental factors, including temperature, humidity, pressure, and pollution, among others. Traditional tools like thermometers and barometers can effectively measure parameters such as temperature and pressure. When it comes to assessing air quality, specialized sensors come into play. These sensors detect the presence of gases and particulate matter in the air [23].
      Additionally, chemical sensors play a crucial role in detecting both chemical and biochemical substances. Among the innovative technologies available are the electronic nose (e-nose) and electronic tongue (e-tongue), which rely on pattern recognition to sense chemicals based on odor and taste. These sensors find valuable applications in smart cities for monitoring pollution levels [24].
@@
      RFID
      RFID technology finds widespread use in various IoT applications, serving as sensors. For instance, it is employed for tracking products within extensive inventories or monitoring items within large retail stores.
@@
    Actuators
    Actuators hold a crucial role and operate in direct contrast to sensors. They transform energy into physical motion and are typically positioned on the outer periphery of a system. Take, for instance, a scenario involving a smart home system that incorporates numerous sensors and actuators. In this setup, the actuators receive signals from the sensors and, depending on the context, carry out actions such as locking or unlocking doors, toggling lights or electrical devices on or off, regulating the house’s temperature, or setting alarms for emergencies. Essentially, actuators respond to and execute commands based on the signals they receive from sensors or other devices.
@@
    Gateway
    The IoT gateway are designed to simplify the IoT system. The gateway functions as the intermediate medium of communication between the devices and the central cloud system. The major functionality of IoT gateways are listed below:
@@
      The IoT gateway acts as an intermediary between sensor devices and the central cloud. Prior to forwarding data directly to the cloud, basic data analytics are performed in this stage. This layer performs the tasks of local data filtering, cleaning, preprocessing, and protocol translation. During this process, it may also aggregate, remove duplicates, or summarize the data to enhance response times and lower transmission costs [25].
@@
      In this layer, data is collected from multiple sources and then transferred to the processing layers. Data is collected and converted into desired format in this layer by utilizing Data acquisition systems. The role of gateway in this stage is to provide secure connectivity between IoT devices and processing structure [26].
@@
      This layer cacilitates real-time device configuration, allowing adjustments to device statuses, operational modes, error acknowledgments, and more [27].
@@
    Cloud
    The cloud serves as the central hub of an IoT network, taking on pivotal roles in data processing, storage, and management. Among its various responsibilities, the following key characteristics stand out:
    1. (a)
      Data Storage and Processing
      The cloud infrastructure serves as the storage solution for the extensive data generated by IoT devices, addressing the impracticality of storing such massive volumes within individual systems. Additionally, data processing takes place within this stage.
    2. (b)
      Scalability
      Cloud infrastructures, including compute, storage, and networking resources, can effortlessly handle thousands of devices.
    3. (c)
      Flexibility
      Cloud-based IoT offers considerable flexibility, allowing devices to be added or removed as needed without requiring a complete system reconfiguration.
    4. (d)
      Maintenance
      In cloud-based IoT, network devices are supervised and managed by the cloud service provider.
    5. (e)
      Cost Effective
      Cloud usage is cost-effective as users only incur charges for the resources they actively utilize.
@@
    Analytics
    This represents one of the most intricate and vital layers within IoT. It involves the analysis of data, generating valuable insights through the application of diverse machine learning algorithms and statistical analysis techniques. Numerous applications of analytics in IoT encompass Anomaly Detection, Environmental Monitoring, Energy Management, Smart Cities, and Agriculture [29].

### 12 Jan 2024 v2

@@ Abstract
The Internet of Things has rapidly transformed the 21st century, enhancing decision-making processes and introducing innovative consumer services such as pay-as-you-use models. The integration of smart devices and automation technologies has revolutionized every aspect of our lives, from health services to the manufacturing industry, and from the agriculture sector to mining. Alongside the positive aspects, it is also essential to recognize the significant safety, security, and trust concerns in this technological landscape. This chapter serves as a comprehensive guide for newcomers interested in the IoT domain, providing a foundation for making future contributions. Specifically, it discusses the overview, historical evolution, key characteristics, advantages, architectures, taxonomy of technologies, and existing applications in major IoT domains. In addressing prevalent issues and challenges in designing and deploying IoT applications, the chapter examines security threats across architectural layers, ethical considerations, user privacy concerns, and trust-related issues. This discussion equips researchers with a solid understanding of diverse IoT aspects, providing a comprehensive understanding of IoT technology along with insights into the extensive potential and impact of this transformative field.
Keywords: Architectural Layers, Ethical Considerations, Security Threats, Hardware Platforms, Trust-related Issues, Taxonomy of technologies.
@@ 1.1 Introduction
The IoT is a network of interconnected devices with sensors, actuators, processors, and various communication technologies. Sensors collect real-time data from both internal states and external surroundings, ranging from mobile phones to microwave ovens. Actuators, in turn, respond to data or commands, enabling automation and remote control of physical devices. Data collected by sensors undergoes processing either at the network edge or on central servers, with some preprocessing occurring directly in the sensors or end devices. Processed data is then transmitted to remote servers for further analysis, storage, and processing. These data form the basis for analysis, decision-making, and subsequent actions, which can be physical (e.g., adjusting a smart thermostat) or virtual (e.g., sending notifications) [4]. The applications of IoT are extensive and diverse, impacting various aspects of our lives. From personal convenience in smart homes to healthcare and fitness innovations, IoT has the potential to influence personal, financial, physical, educational, professional, and mental aspects of individuals. In smart homes, IoT enables remote control of electrical appliances, lighting, coffee brewing, thermostat adjustments, and even hands-free operations through voice commands [5]. In healthcare, wearable IoT devices offer remote monitoring, allowing caregivers and healthcare professionals to provide timely assistance in emergencies. Additionally, individuals can use wearable devices to track sleep patterns, physical activity, and overall fitness [6]. These examples only scratch the surface of IoT’s broad application landscape, indicating the exciting possibilities and challenges that researchers are exploring for the future.
@@
The IoT has the potential to transform how people interact with technology, providing greater convenience, efficiency, and personalization in daily life. Despite its transformative impact, IoT faces challenges. The sheer number of devices and the substantial data generated pose significant challenges, with a projected 41.6 billion IoT devices producing 79.4 zettabytes of data by 2025 [7]. Addressing this requires scalable architectures and enhanced processing capabilities. Moreover, IoT heavily relies on wireless communication, leading to challenges such as distortion and unreliability in geographically dispersed locations. Ensuring dependable data transmission becomes a pivotal challenge, emphasizing the critical role of communication technologies in the IoT landscape. Beyond technical hurdles, various general and domain-specific challenges are crucial for the success of IoT. Identifying and addressing these multifaceted challenges collectively is essential to unlocking the full potential of IoT and overcoming obstacles for broader adoption and integration into our lives.
@@
Motivation of this chapter: IoT is not a novel concept; researchers have been exploring this field for decades. Consequently, the question arises: what sets this chapter apart from others, and why should readers invest their time in exploring the basics presented here? While there are several book chapters introducing IoT, many share outdated concepts, lacking the latest insights from the ongoing exploration of this dynamic field. For instance, in [8], only IoT applications are discussed, and in [9], authors overlook major domains, including IoT applications, advantages, challenges, and technologies. Notably, Nagaraj et al. [10] present a well-structured discussion covering technologies, architectures, applications, and challenges. However, their discussion is limited to these aspects, with a less extensive exploration of the application section. Therefore, the need arises for a comprehensive chapter that covers all aspects and incorporates recent additions in the IoT field. This chapter goes beyond the basics, encompassing fundamental components, characteristics, and advantages. It delves into architectures, provides a taxonomy of technologies used in IoT, explores a significant number of applications across diverse domains, and addresses ethical considerations as well as legal and regulatory issues. In essence, our chapter serves as a holistic guide, covering the most significant facets of IoT, from foundational principles to emerging research challenges and future directions. Readers are encouraged to explore this chapter for a thorough and up-to-date understanding of the evolving landscape of IoT.
Contributions: The contributions of this chapter can be summarized as follows:
@@
Table 1.1: List of abbreviations
Name Abbreviation Name Abbreviation Name Abbreviation
FDMA Frequency Division Multiple Access LBT Listen Before Talk CDMA Code Division Multiple Access
FHSS Frequency Hopping Spread Spectrum CSS Chirp Spread Spectrum OOK On-Off Keying
TDMA Time Division Multiple Access GSM
Global System for Mobile
Communications
PPM Pulse Position Modulation
GFSK Gaussian Frequency Shift Keying UMTS
Universal Mobile Telecommunications
System
OFDM
Orthogonal Frequency Division
Multiplexing
DWPSK
Differential Quadrature Phase Shift
Keying
GMSK Gaussian Minimum Shift Keying TDD Time Division Duplex
CSMA/CA
Carrier Sense Multiple Access with
Collision Avoidance
LTE-A Long-Term Evolution Advanced FDD Frequency Division Duplex
BPSK Binary Phase Shift Keying 3GPP 3rd Generation Partnership Project ALHOA Adaptive Link Hopping On Air
CSMA/CD
Carrier Sense Multiple Access with
Collision Detection
WCDMA
Wideband Code Division Multiple
Access
DBPSK
Differential Binary Phase Shift
Keying
O-QPSK Offset Quadrature Phase Shift Keying OFDMA
Orthogonal Frequency Division
Multiple Access
QAM Quadrature Amplitude Modulation
QPSK Quadrature Phase Shift Keying SC-FDMA
Single Carrier Frequency Division
Multiple Access
CP-OFDM
Cyclic Prefix Orthogonal Frequency
Division Multiplexing
Chapter organization: It begins with an exploration of the fundamentals in Section 1.2, covering the history 1.2.1, components 1.2.2, and characteristics 1.2.3. In Section 1.3, the advantages of IoT are examined. By introducing a generic architecture in Section 1.4, the taxonomy of technologies that underpin IoT operations are discussed in detail in Section 1.5. Furthermore, Section 1.6 explores the extensive array of application domains of IoT, with a focus on multiple promising research works within each domain, aiming to provide a comprehensive understanding of its real-world applications. It concludes in Section 1.7 by addressing the key challenges that IoT faces today, considering insights from the reviewed literature, and sheds light on potential future scopes and developments in the ever-evolving realm of IoT technology. Table 1.1 represents some abbreviations mentioned in this chapter.
@@ 1.2 Fundamentals of IoT
At its core, the IoT revolves around the concept of connecting everyday objects and devices to the internet, enabling them to communicate, collect data, and perform actions autonomously or in response to commands. This basic concept involves three key elements: sensors and actuators that gather and interact with data, a network infrastructure for data transmission, and cloud-based platforms for data storage, processing, and analysis. By interconnecting these elements, IoT empowers us to enhance efficiency, gain real-time insights, and create smart, responsive systems that impact various aspects of our lives, from smart homes and cities to industries and healthcare. The evolution of IoT, which has unfolded over several decades, began with the concept of smart objects. Since then, IoT has undergone numerous groundbreaking transformations that have left the world astounded, thanks to its unique and convenient characteristics. The advantages brought by this technology are unquantifiable, as it touches nearly every aspect of people’s lives, simplifying and improving them. This section comprehensively discusses these fundamental IoT concepts, covering the following topics: history, components, characteristics.
1.2.1 Historical Development
@@
In the 1980s, there was a growing interest in enhancing physical objects with sensors and intelligence. Commercial service providers began supporting public access to ARPANET, an early precursor to the modern Internet, during this time. Satellites and landlines played pivotal roles in establishing the foundational communication infrastructure for the emerging IoT. The concept of a network of smart devices was initially explored as early as 1982 when a Coca-Cola vending machine [^3] The ”Only” Coke Machine on the Internet. Available at: Carnegie Mellon University at Carnegie Mellon University was modified to connect to ARPANET. This allowed local programmers to remotely monitor the vending machine’s contents, ensuring drinks were available and cold before making a purchase. However, the technology was challenging to manage, and progress in this field was limited during that period [13]. In parallel, during the 1980s, local area networks (LANs) gained popularity and proved effective for real-time communication and document sharing among groups of PCs.
@@
In the 1990s, advancements in wireless technology set the stage for the widespread adoption of ”machine-to-machine” (M2M) solutions in enterprise and industrial contexts, particularly for equipment monitoring and operation. However, many of these early M2M solutions relied on closed, purpose-built networks and proprietary or industry-specific standards rather than utilizing Internet Protocol (IP)-based networks and Internet standards [14]. By the mid-1990s, the Internet had expanded its global reach, offering new possibilities for researchers and technologists to explore ways to enhance connections between humans and machines. A significant milestone in this journey was the creation of the first Internet-connected ’device’ by John Romkey—an IP-enabled toaster that could be controlled over the Internet. This innovative toaster was showcased at an Internet conference in 1990, marking an early example of the IoT in action [15].
@@
In 1991, Mark Weiser’s paper “The Computer of the 21st Century” [16] and academic events like UbiComp and PerCom shaped the contemporary IoT vision [14]. Global Positioning System (GPS) became a reality in early 1993 with the Department of Defense establishing a stable system of 24 satellites. Privately owned commercial satellites soon followed, enhancing IIoT functionality [3]. In early 1994, Reza Raji introduced the IoT concept in IEEE Spectrum, describing it as “moving small data packets to integrate and automate from home appliances to entire factories” [17]. Later that year, Steve Mann invented the near-real-time WearCam, powered by a 64-processor setup. Between 1993 and 1997, companies proposed IoT solutions, including Microsoft’s “at Work” and Novell’s NEST. Momentum grew as Bill Joy introduced device-to-device communication in his ”Six Web” framework at the 1999 World Economic Forum [3].
@@
The term “Internet of Things” was coined by Peter T. Lewis in a 1985 speech during the Congressional Black Caucus Foundation’s 15th Annual Legislative Weekend in Washington, D.C. Lewis defined IoT as the integration of people, processes, and technology with connectable devices and sensors for remote monitoring, status assessment, manipulation, and trend evaluation related to these devices [3]. In 1997, Paul Saffo described sensors and their future roles. British technologist Kevin Ashton, while serving as the executive director of the Auto-ID Center at MIT, independently coined the term “Internet of Things”. During his time at Procter and Gamble, he explored radio-frequency identification (RFID), a technology framework enabling physical devices to connect via microchips and wireless signals. In the same year, they developed a global RFID-based item identification system [14].
@@
In 1999, Kevin Ashton was the first to describe the IoT and proposed the name “Internet of Things” during a presentation for Procter and Gamble. He believed RFID technology, primarily designed for inventory tracking, was a significant prerequisite for the IoT, allowing computers to efficiently manage and monitor individual objects. The concept of tagging objects has been realized through technologies like digital watermarking, barcodes, and QR codes, used for identification and tracking purposes [18]. Subsequent technological advancements, including the proliferation of smartphones, cloud computing, improved processing power, and enhanced software algorithms, along with the availability of sophisticated sensors capable of measuring various parameters, laid the foundation for robust data collection, storage, and processing for the IoT’s growth.
@@
As a significant step forward in commercializing IoT, LG announced plans to launch a smart refrigerator capable of autonomously managing its contents in 2000. Walmart and the US Department of Defense pioneered inventory tracking using RFID and the IoT in $2002-2003$. RFID gained prominence in the US Army’s Savi program in 2003, and Walmart expanded its RFID usage worldwide that same year. In 2004, Cornelius “Pete” Peterson, CEO of NetSilicon, predicted that IoT devices would dominate the next era of information technology, particularly in fields like medical devices and industrial controls [3]. In 2005, numerous articles in mainstream newspapers such as The Guardian, Scientific American, and The Boston Globe discussed IoT’s future direction.
@@
1.2.2 Components
The IoT consists of several key components that serve as the essential building blocks for constructing an IoT system. This section provides an in-depth exploration of the principal components of the IoT. IoT comprises three main components: (1) sensors/devices, and actuators; (2) storage and data analytics; and (3) interpretation and visualization tools. Each of these is further categorized into various subcomponents.
@@ 1.2.2 Components
    Sensors/Devices: Sensors play a crucial and essential role within an IoT system. Given that IoT operates by gathering data from the surrounding environment, it is necessary for all IoT applications to incorporate one or more sensors to meet this need. A defining characteristic of IoT devices is their context awareness, which is made possible through the utilization of sensor technology. Sensors are not only compact and cost-effective but also energy-efficient. However, they are subject to limitations such as battery capacity and ease of deployment. [4]. An overview of various types of sensors has been provided below.
@@
      Mobile-based Sensors
      Smartphones, which are widespread and commonly used, are equipped with various sensors. Given their extensive usage, researchers are exploring the potential of using smartphones as integral components in building smart IoT solutions. These applications can harness sensor data from smartphones to generate valuable insights and outcomes. Some of the general sensors found in smartphones include an accelerometer, gyroscope, GPS, magnetometer, light sensor, and proximity sensor [19]. Certain smartphones, like the Samsung Galaxy S4, come equipped with extra sensors, including a thermometer, barometer, and humidity sensor [4].
@@
      The healthcare industry is one of the most influential fields where innovation and IoT have paved the way. Wearable devices and sensors have facilitated remote monitoring for physicians and enabled researchers to collect data continuously and in real-time. These devices come in different forms, such as wristbands, smartwatches, and monitoring patches. Smartwatches and fitness trackers, known for their versatility, have gained popularity among consumers. Likewise, monitoring patches have emerged as a valuable asset to the healthcare sector by enabling remote treatment for patients.
@@
      Neural sensors play a crucial role in comprehending the workings of human neurons by enabling us to decode brain signals, assess the brain’s current state, and, when necessary, optimize it for improved focus and attention. This practice is commonly referred to as Neurofeedback [22].
@@
      While conventional tools manage parameters such as temperature and pressure, specialized environmental sensors play a crucial role in evaluating air quality. These sensors detect gases and particulate matter [23], while also measuring factors like temperature, humidity, pressure, and pollution. Besides, chemical sensors play a crucial role in detecting both chemical and biochemical substances. Among the innovative technologies available are the electronic nose (e-nose) and electronic tongue (e-tongue), which rely on pattern recognition to sense chemicals based on odor and taste. These sensors find valuable applications in smart cities for monitoring pollution levels [24].
@@
      Radio-Frequency Identification (RFID)
      RFID technology, which serves as sensors, finds widespread use in various IoT applications. For instance, it is employed for tracking products within extensive inventories or monitoring items within large retail stores.
@@
    Actuators: Actuators hold a crucial role and operate in direct contrast to sensors. They transform energy into physical motion and are typically positioned on the outer periphery of a system. Take, for instance, a scenario involving a smart home system that incorporates numerous sensors and actuators. In this setup, the actuators receive signals from the sensors and, depending on the context, carry out actions such as locking or unlocking doors, toggling lights or electrical devices on or off, regulating the house’s temperature, or setting alarms for emergencies. Essentially, actuators respond to and execute commands based on the signals they receive from sensors or other devices.
@@
    Gateway: Gateways, designed to simplify the IoT system, function as the intermediate medium of communication between the devices and the central cloud system. The major functionalities of IoT gateways are listed below:
@@
      The IoT gateway serves as an intermediary between sensor devices and the central cloud, conducting basic data analytics before forwarding information directly to the cloud. This layer performs the tasks of local data filtering, cleaning, preprocessing, and protocol translation. During this process, it may also aggregate, remove duplicates, or summarize the data to enhance response times and lower transmission costs [25].
@@
      In this layer, data is collected from multiple sources, converted into the desired format, and then transferred to the processing layers. The role of the gateway in this stage is to provide secure connectivity between IoT devices and processing structures [26].
@@
      This layer facilitates real-time device configuration, allowing adjustments to device statuses, operational modes, error acknowledgments, and more [27].
@@
    Cloud: The cloud serves as the central hub of an IoT network, taking on pivotal roles in data processing, storage, and management. The key characteristics of the cloud include the ability to store and process extensive data generated by devices, scalability to effortlessly handle thousands of devices, flexibility by allowing devices to be added or removed as needed without requiring a complete system reconfiguration, supervision and management by the cloud service provider, and cost-effectiveness.
@@
    Analytics: This represents one of the most intricate and vital layers within IoT. It involves the analysis of data, generating valuable insights through the application of diverse machine learning (ML) algorithms and statistical analysis techniques. Numerous applications of analytics in IoT encompass anomaly detection, environmental monitoring, energy management, smart cities, and agriculture [29].
Figure 1.1: IoT characteristics